# The Nontrivial Vacuum Structure of an Extended $t\overline{t}$ BEH (Higgs) Bound state

### v1

@@ III.2 Discrete Summation
$S_{3}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(2{g^{\prime}{}_{0}^{2}N_{c}}D_{F}(2r^{\mu})|H^{\prime}|^{2}|\Phi(r))|^{2}\biggr)\approx M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;{|H^{\prime}|^{2}}\biggl(2g^{\prime}{}_{0}^{2}N_{c}\frac{\delta^{4}(r^{\mu})}{16M_{0}^{2}}\frac{1}{N}\sum^{N}_{i=1}\phi_{\omega_{i}}^{\dagger}(r)\sum^{N}_{j=1}\phi_{\omega_{j}}(r)\biggr)$
$=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;N|H^{\prime}|^{2}\biggl(\frac{g^{\prime}{}_{0}^{2}N_{c}}{8M_{0}^{2}}\delta^{4}(r)\phi^{\dagger}_{\omega}(0)\phi_{\omega}(0)\biggr)=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;|H|^{2}\biggl(\frac{g^{\prime}{}_{0}^{2}N_{c}}{8M_{0}^{2}}\delta^{4}(r)\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r)\biggr)\;(\text{for any }\omega)$ (43)
The essential result is that the potential is approximately $\sim\delta^{4}(r^{\mu})$, hence the line integral orientation $\omega_{\mu}$ becomes irrelevant in the potential. In the second-to-last term we see the usual enhancement factor, $N$, that would normally lead to the BEC or BCS phenomena. However, in the last term we see that the normalization of $H^{\prime}=H/\sqrt{N}$ undoes the $N$-fold enhancement and the coupling is not renormalized $g_{0}^{\prime 2}=g_{0}^{2}$. This is therefore different than the case of the BCS superconductor or the BEC where the coupling is enhanced by a factor of $N$. The nonrenormalization of $g_{0}^{2}$ owes to the bilocal field theory with the renormalization of $H^{\prime}=H/\sqrt{N}$.
@@
$S_{3}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r|H(X)|^{2}\biggl(\frac{2g_{0}^{2}N_{c}}{M_{0}^{2}}D_{F}(2r)|\widetilde{\Phi}(r)|^{2}\biggr).$ (44)
To a good approximation, we can freely swap between $D_{F}(2r)\leftrightarrow\delta^{4}(r)/16M_{0}^{2}$.
@@
$=g_{Y}M_{0}^{2}\int\!\!d^{4}X\;[\overline{\psi}_{iL}(X)\psi_{R}(X)]_{f}\;\;H^{i}(X)\widetilde{\Phi}(0){+h.c.};\qquad\qquad g_{Y}=g_{0}^{2}\sqrt{2N_{c}/J}$ (45)
where we have taken the limit $D_{F}(2r)\rightarrow\delta^{4}(r)/16M_{0}^{2}$. (Here $J=16$ is the Jacobian in passing from coordinates $(x,y)$ to $(X,r)$ which is derived in chris1 )
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{1}{2}\hat{\lambda}^{\prime}\int d^{4}Xd^{4}r\;\delta^{4}(r)|H^{\prime}(X)\Phi(r)|^{4}=\frac{1}{2}\hat{\lambda}\int d^{4}X|H(X)|^{4}\left|\widetilde{\Phi}(0)\right|^{4}$ (46)
where $\lambda^{\prime}=\lambda$ is therefere not renormalized. Essentally $\lambda$ is determined by the Yukawa coupling, $g_{Y}$ and the renormalization group running $\lambda$ from $M_{0}$ to $|\mu|$ and the result we obtained from the underlying theory at one loop, e.g., $\lambda\approx 0.23$, is not modifed. This term would also permit expansion in $r^{\mu}$ and yield an operator expansion of new physics.
@@
$S_{\Phi}=Z^{\prime}M_{0}^{4}\mu^{2}\int d^{4}Xd^{4}r|H^{\prime}|^{2}\;|\widetilde{\Phi}|^{2}=\mu^{2}\!\!\int d^{4}X|H|^{2}$ (47)
We thus obtain the same eigenvalue for $\Phi(r)$ as for the underlying $\phi_{\omega}(r)$.
III.3 Continuous Integral Representation
@@
$\sum_{i}^{N}(c)={\cal{N^{\prime}}}\int d^{4}\omega\;\delta(\omega^{2}-1)\;\phi_{\omega}(r^{\mu})\equiv\int_{\omega}(c)=N(c)$ (48)
@@
$\Phi(r^{\mu})=\frac{1}{\sqrt{N}}\sum_{i}^{N}\phi_{\omega_{i}}={\cal{N}}\int d^{4}\omega\;\delta(\omega^{2}-1)\;\phi_{\omega}(r^{\mu})\equiv\frac{1}{\sqrt{N}}\int_{\omega}\phi_{\omega}(r^{\mu})$ (49)
@@
Note that $\cal{N}$ is the normalization of a divergent integral over the hyperboloid, and requires, in principle, regularization. We will not enter into a detailed discussion of regularization here. We note, however, that we can usually perform a “Wick rotation” if the integral is an analytic function of the metric, $g_{\mu\nu}=(1,-1,-1,-1)$ it can be continued as $g_{\mu\nu}\rightarrow\eta_{\mu\nu}\sim(1,1,1,1)$. Then the hyperboloid is replaced by a Euclidean 4-sphere and
@@
$Z^{\prime}M_{0}^{4}\int d^{4}Xd^{4}r|H^{\prime}|^{2}|\Phi(r)|^{2}=Z^{\prime}M_{0}^{4}\int d^{4}Xd^{4}r|H^{\prime}|^{2}\frac{1}{\sqrt{N}}\int_{\omega}\phi^{\dagger}_{\omega}(r)\frac{1}{\sqrt{N}}\int_{\omega^{\prime}}\phi^{\dagger}_{\omega^{\prime}}(r)$
$=ZM_{0}^{4}\int d^{4}Xd^{4}r|H|^{2}\frac{1}{{N}}\int_{\omega}\phi^{\dagger}_{\omega}(r)\phi^{\dagger}_{\omega}(r)\times\biggl(\int_{\omega^{\prime}}\biggr)=ZM_{0}^{4}\int d^{4}Xd^{4}r|H|^{2}$ (51)
@@
$\int\!\!d^{4}r\;\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{j}}(r)=\delta_{ij}\int\!\!d^{4}r\;\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r)=\delta_{ij}\!\!\int d^{4}r|\widetilde{\Phi}(r^{\mu})|^{2}$ (52)
@@ IV Full Action
The full action then becomes that of the standard model BEH boson coupled to top quarks with in the unprimed parameters:
$S=\!\!\int\!\!d^{4}X\;\biggl(|DH(X)|^{2}-\mu^{2}|H|^{2}+g_{Y}H^{i\dagger}(X)[\overline{\psi}_{R}(X)\psi_{iL}(X)+h.c.]_{f}-\frac{1}{2}\lambda(H^{\dagger}H)^{2}\biggr)+S^{\prime},$ (54)
@@
$\mu^{2}|H|^{2}+\frac{1}{2}\lambda(H^{\dagger}H)^{2}\qquad\text{where,}\;\;\;\mu^{2}<0$ (55)
@@
$H^{\prime}(X){\Phi}(r)=H(X)\widetilde{\Phi}(r)\rightarrow\exp(i\pi^{a}(X)\tau^{a}/2v_{weak})\biggl(\begin{array}[]{c}v_{weak}\widetilde{\Phi}(r^{\mu})+\frac{h(X,r^{\mu})}{\sqrt{2}}\\ 0\end{array}\biggr)$ (58)
where $h(X,r^{\mu})=h(X)\widetilde{\Phi}(r^{\mu})$ is the physical BEH boson and $\pi^{a}(X)$ are Nambu-Goldstone bosons, which are phase factors of $H(X)$. The electroweak symmetry is broken spontaneously and the gauge fields absorb the Nambu-Goldstone phase factors and acquire mass in the usual way. The resulting broken phase of the BEH field, $h(X)$ is then the standard model and the canonical normalization of $h(X)$ follows from the $Z$ normalization of the internal field $\widetilde{\Phi}(r)$ in eq.(28).
@@
$H^{\prime}(X)\Phi=H(X)\widetilde{\Phi}(r)\rightarrow v_{weak}\widetilde{\Phi}(r)+h(X)\widetilde{\Phi}(r)/\sqrt{2}$ (59)
where $h(X)$ is the SM BEH boson. One might be concerned that the “two-body” field, $h(X)$, is now associated with apparent relative time through $\widetilde{\Phi}(r)$, however, $\widetilde{\Phi}(r)$ is a vacuum feature and makes no reference to a particular $\omega_{\mu}$. Hence the SM BEH field, $h(X)$ is a collective object and its precise two-body nature is actually blended wth the collective vacuum $\Phi$. Likewise, the Nambu-Goldstone bosons are “eaten” by the gauge fields and $\Phi(r)$ is a spectator to the usual combinatorics of that.
We obtain the BEH kinetic term eq.III.2) of the SM :
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{1}\rightarrow\frac{1}{2}ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;(\partial h(X))^{2}|\widetilde{\Phi}(r)|^{2}=\frac{1}{2}\int\!\!d^{4}X(\partial h)^{2}$ (60)
where we integrate out $\widetilde{\Phi}$ in eqs.(III.2,60) using eq.(41). In fact, there is no way to discern the compositeness of the BEH field from the kinetic and mass terms (this requires the Yukawa interaction and quartic terms). Moreover, $\langle H\rangle\rightarrow v_{weak}$ with the covariant derivative, $D_{\mu}$, of eq.(74), leads to:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{1}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z|DH(X)|^{2}|\widetilde{\Phi}(r)|^{2}\biggr)\rightarrow$
$=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(M_{W}^{2}W^{+}W^{-}+\frac{1}{2}M_{Z}Z^{2}\biggr)|\widetilde{\Phi}(r)|^{2}=\!\int\!\!d^{4}X\biggl(M_{W}^{2}W^{+}W^{-}+\frac{1}{2}M_{Z}Z^{2}\biggr)$ (61)
@@
$\int d^{4}X\biggl(\frac{1}{2}\partial_{\mu}h\partial^{\mu}h-\frac{1}{2}m_{h}^{2}h^{2}-\sqrt{\frac{\lambda}{2}}|\mu|h^{3}-\frac{\lambda}{8}h^{4}\biggr)+\text{electroweak couplings}$ (62)
@@
The structure of the collective field, $\Phi(r)$, can in principle be probed through the extended interactions, but the effects will be suppressed. The Yukawa coupling term in natural top condensation takes the form chris1 ; chris2 in the broken phase:
$\!\!\!\!\!\!\!\!\!\!S_{Y}^{\prime}=\sqrt{2N_{c}J}g^{2}_{0}M_{0}^{2}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{L}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}D_{F}(2r^{\mu})\;\biggl(v_{weak}+\frac{h(X)}{\sqrt{2}}\biggr)\widetilde{\Phi}(r^{\mu}){+h.c.}.$ (63)
where $\psi_{L}=(t,b)_{L}$ and $\psi_{R}=t_{R}$. $J=16$ is a Jacobian. To a leading approximation we can take the pointlike limit of the potential, and we have $D_{F}(2r^{\mu})\rightarrow(JM_{0}^{2})^{-1}\delta^{4}(r)$. Hence
$g_{Y}=\hat{g}_{Y}|\widetilde{\Phi}(0)|\qquad\makebox{where,}\qquad\hat{g}_{Y}\equiv g_{0}^{2}\sqrt{2N_{c}/J}=g_{0}^{2}\sqrt{3/8}.$ (64)
@@
$\!\!\!\!\!\!\!\!\!\!S_{Y}^{\prime}=g_{Y}\int\!\!d^{4}X\;[\overline{\psi}_{L}(X)\psi_{R}(X]_{f}\biggl(v_{weak}+\frac{h(X)}{\sqrt{2}}\biggr){+h.c.}$ (65)

### v3

@@ III.1 Formal Derivation
$S_{3}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(2{g^{\prime}{}_{0}^{2}N_{c}}D_{F}(2r^{\mu})|H^{\prime}|^{2}|\Phi(r^{\mu}))|^{2}\biggr)\approx M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;{|H^{\prime}|^{2}}\biggl(2g^{\prime}{}_{0}^{2}N_{c}\frac{\delta^{4}(r^{\mu})}{16M_{0}^{2}}\frac{1}{N}\sum^{N}_{i=1}\phi_{\omega_{i}}^{\dagger}(r)\sum^{N}_{j=1}\phi_{\omega_{j}}(r)\biggr)$
$=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;N|H^{\prime}|^{2}\biggl(\frac{g^{\prime}{}_{0}^{2}N_{c}}{8M_{0}^{2}}\delta^{4}(r)\phi^{\dagger}_{\omega}(0)\phi_{\omega}(0)\biggr)=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;|H|^{2}\biggl(\frac{g^{\prime}{}_{0}^{2}N_{c}}{8M_{0}^{2}}\delta^{4}(r)\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r)\biggr)\;(\text{for any }\omega).$ (42)
The essential result is that the potential is approximately $\sim\delta^{4}(r^{\mu})$, hence the line integral orientation $\omega^{\mu}$ becomes irrelevant in the potential. In the second-to-last term we see the usual enhancement factor, $N$, that would normally lead to the BEC or BCS phenomena. However, in the last term we see that the normalization of $H^{\prime}=H/\sqrt{N}$ undoes the $N$-fold enhancement and the coupling is not renormalized $g_{0}^{\prime 2}=g_{0}^{2}$. This is therefore different than the case of the BCS superconductor or the BEC where the coupling is enhanced by a factor of $N$. The nonrenormalization of $g_{0}^{2}$ owes to the bilocal field theory with the renormalization of $H^{\prime}=H/\sqrt{N}$.
@@
$S_{3}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r|H(X^{\mu})|^{2}\biggl(\frac{2g_{0}^{2}N_{c}}{M_{0}^{2}}D_{F}(2r)|\widetilde{\Phi}(r)|^{2}\biggr).$ (43)
To a good approximation we can freely swap between $D_{F}(2r)\leftrightarrow\delta^{4}(r)/16M_{0}^{2}$.
@@
$=g_{Y}M_{0}^{2}\int\!\!d^{4}X\;[\overline{\psi}_{iL}(X)\psi_{R}(X)]_{f}\;\;H^{i}(X)\widetilde{\Phi}(0){+h.c.};\qquad\qquad g_{Y}=g_{0}^{2}\sqrt{2N_{c}/J},$ (44)
where we have taken the limit $D_{F}(2r)\rightarrow\delta^{4}(r)/16M_{0}^{2}$.
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{1}{2}\hat{\lambda}^{\prime}\int d^{4}Xd^{4}r\;\delta^{4}(r)|H^{\prime}(X)\Phi(r^{\mu})|^{4}=\frac{1}{2}\hat{\lambda}\int d^{4}X|H(X^{\mu})|^{4}|\widetilde{\Phi}(0)|^{4},$ (45)
where $\lambda^{\prime}=\lambda$ is therefore not renormalized. Essentially $\lambda$ is determined by the Yukawa coupling, $g_{Y}$ and the renormalization group running $\lambda$ from $M_{0}$ to $|\mu|$ and the result we obtained from the underlying theory at one loop, e.g., $\lambda\approx 0.23$, is not modified. This term would also permit expansion in $r^{\mu}$ and yield an operator expansion of new physics.
@@
$S_{\Phi}=Z^{\prime}M_{0}^{4}\mu^{2}\int d^{4}Xd^{4}r|H^{\prime}|^{2}\;|\widetilde{\Phi}|^{2}=\mu^{2}\!\!\int d^{4}X|H|^{2}.$ (46)
We thus obtain the same eigenvalue for $\Phi(r^{\mu})$ as for the underlying $\phi_{\omega}(r)$.
@@
$\sum_{i}^{N}(c)={\cal{N^{\prime}}}\int d^{4}\omega\;\delta(\omega^{2}-1)\;\phi_{\omega}(r^{\mu})\equiv\int_{\omega}(c)=N(c),$ (47)
@@
$\Phi(r^{\mu})=\frac{1}{\sqrt{N}}\sum_{i}^{N}\phi_{\omega_{i}}={\cal{N}}\int d^{4}\omega\;\delta(\omega^{2}-1)\;\phi_{\omega}(r^{\mu})\equiv\frac{1}{\sqrt{N}}\int_{\omega}\phi_{\omega}(r^{\mu}).$ (48)
@@
Note that $\cal{N}$ is the normalization of a divergent integral over the hyperboloid, and requires, in principle, regularization. We will not enter into a detailed discussion of regularization here, however we note that we can usually perform a “Wick rotation.” If the integral is an analytic function of the metric, $g_{\mu\nu}=(1,-1,-1,-1)$, it can be continued as $g_{\mu\nu}\rightarrow\eta_{\mu\nu}\sim(1,1,1,1)$. Then the hyperboloid is replaced by a Euclidean 4-sphere and
@@
$Z^{\prime}M_{0}^{4}\int d^{4}Xd^{4}r|H^{\prime}|^{2}|\Phi(r^{\mu})|^{2}=Z^{\prime}M_{0}^{4}\int d^{4}Xd^{4}r|H^{\prime}|^{2}\frac{1}{\sqrt{N}}\int_{\omega}\phi^{\dagger}_{\omega}(r)\frac{1}{\sqrt{N}}\int_{\omega^{\prime}}\phi^{\dagger}_{\omega^{\prime}}(r)$
$=ZM_{0}^{4}\int d^{4}Xd^{4}r|H|^{2}\frac{1}{{N}}\int_{\omega}\phi^{\dagger}_{\omega}(r)\phi^{\dagger}_{\omega}(r)\times\biggl(\int_{\omega^{\prime}}\biggr)=ZM_{0}^{4}\int d^{4}Xd^{4}r|H|^{2},$ (50)
@@
$\int\!\!d^{4}r\;\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{j}}(r)=\delta_{ij}\int\!\!d^{4}r\;\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r)=\delta_{ij}\!\!\int d^{4}r|\widetilde{\Phi}(r^{\mu})|^{2},$ (51)
@@ IV Full Action
The full action in the pointlike limit then becomes that of the standard model BEH boson coupled to top quarks with in the unprimed parameters:
$S=\!\!\int\!\!d^{4}X\;\biggl(|DH(X^{\mu})|^{2}-\mu^{2}|H|^{2}+g_{Y}H^{i\dagger}(X)[\overline{\psi}_{R}(X)\psi_{iL}(X)+h.c.]_{f}-\frac{1}{2}\lambda(H^{\dagger}H)^{2}\biggr)+S^{\prime},$ (53)
@@ IV.1 BEH Potential
$\mu^{2}|H|^{2}+\frac{1}{2}\lambda(H^{\dagger}H)^{2},\qquad\text{where,}\;\;\;\mu^{2}<0.$ (54)
@@
$H^{\prime}(X){\Phi}(r)=H(X^{\mu})\widetilde{\Phi}(r)\rightarrow\exp(i\pi^{a}(X)\tau^{a}/2v_{weak})\biggl(\begin{array}[]{c}v_{weak}\widetilde{\Phi}(r^{\mu})+\frac{h(X,r^{\mu})}{\sqrt{2}}\\ 0\end{array}\biggr),$ (57)
where $h(X,r^{\mu})=h(X)\widetilde{\Phi}(r^{\mu})$ is the physical BEH boson and $\pi^{a}(X)$ are Nambu-Goldstone bosons, which are phase factors of $H(X^{\mu})$. The electroweak symmetry is broken spontaneously and the gauge fields absorb the Nambu-Goldstone phase factors and acquire mass in the usual way. The resulting broken phase of the BEH field, $h(X)$ is then the standard model and the canonical normalization of $h(X)$ follows from the $Z$ normalization of the internal field $\widetilde{\Phi}(r)$ in eq.(27).
@@
$H^{\prime}(X)\Phi=H(X^{\mu})\widetilde{\Phi}(r)\rightarrow v_{weak}\widetilde{\Phi}(r)+h(X)\widetilde{\Phi}(r)/\sqrt{2},$ (58)
where $h(X)$ is the SM BEH boson. The “two-body” field, $h(X)$, is now associated with apparent relative time through $\widetilde{\Phi}(r)$, however, the key feature is that $\widetilde{\Phi}(r)\sim\widetilde{\Phi}(0)+{\cal{O}}(r_{\mu}r^{\mu})+$... is a Lorentz invariant function of $r_{\mu}r^{\mu}$ and makes no reference to a particular $\omega^{\mu}$. This expansion can, in principle, yield a form factor in BEH interactions and/or higher dimension operators as described above. Hence the SM BEH field, $h(X)$ is a collective object and its precise two-body nature is actually blended with the collective vacuum $\Phi$. Likewise, the Nambu-Goldstone bosons are “eaten” by the gauge fields in the usual way, and $\Phi(r^{\mu})$ is integrated out and does not affect the usual combinatorial.
The presence of the collective wave-function $\Phi(r^{\mu})$ is not detectable in the kinetic terms, nor in the interaction in the pointlike limit. We obtain the BEH kinetic term eq.III.1) of the SM :
$+\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{1}\rightarrow\frac{1}{2}ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;(\partial h(X))^{2}|\widetilde{\Phi}(r)|^{2}=\frac{1}{2}\int\!\!d^{4}X(\partial h)^{2},$ (59)
where we integrate out $\widetilde{\Phi}$ in eqs.(III.1,59) using eq.(40). In fact, there is no way to discern the compositeness of the BEH field from the kinetic and mass terms (this requires the Yukawa interaction and quartic terms). Moreover, $\langle H\rangle\rightarrow v_{weak}$ with the covariant derivative, $D_{\mu}$, of eq.(80), leads to:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{1}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z|DH(X^{\mu})|^{2}|\widetilde{\Phi}(r)|^{2}\biggr)\rightarrow$
$=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(M_{W}^{2}W^{+}W^{-}+\frac{1}{2}M_{Z}Z^{2}\biggr)|\widetilde{\Phi}(r)|^{2}=\!\int\!\!d^{4}X\biggl(M_{W}^{2}W^{+}W^{-}+\frac{1}{2}M_{Z}Z^{2}\biggr).$ (60)
@@
$\int d^{4}X\biggl(\frac{1}{2}\partial_{\mu}h\partial^{\mu}h-\frac{1}{2}m_{h}^{2}h^{2}-\sqrt{\frac{\lambda}{2}}|\mu|h^{3}-\frac{\lambda}{8}h^{4}\biggr)+\text{electroweak couplings},$ (61)
@@
The structure of the collective field, $\Phi(r^{\mu})$, can in principle be probed through the extended interactions, but the effects will be suppressed. The Yukawa coupling term in natural top condensation takes the form chris1 ; chris2 in the broken phase:
$\!\!\!\!\!\!\!\!\!\!S_{Y}^{\prime}=\sqrt{2N_{c}J}g^{2}_{0}M_{0}^{2}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{L}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}D_{F}(2r^{\mu})\;\biggl(v_{weak}+\frac{h(X)}{\sqrt{2}}\biggr)\widetilde{\Phi}(r^{\mu}){+h.c.},$ (62)
where $\psi_{L}=(t,b)_{L}$ and $\psi_{R}=t_{R}$ (and $J=16$ is the Jacobian in passing from coordinates $(x,y)$ to $(X,r)$ chris1 ).
To a good approximation we can take the pointlike limit of the potential, and we have $D_{F}(2r^{\mu})\rightarrow(JM_{0}^{2})^{-1}\delta^{4}(r)$. Hence
$g_{Y}=\hat{g}_{Y}|\widetilde{\Phi}(0)|,\qquad\makebox{where,}\qquad\hat{g}_{Y}\equiv g_{0}^{2}\sqrt{2N_{c}/J}=g_{0}^{2}\sqrt{3/8}.$ (63)
@@
$\!\!\!\!\!\!\!\!\!\!S_{Y}^{\prime}=g_{Y}\int\!\!d^{4}X\;[\overline{\psi}_{L}(X)\psi_{R}(X]_{f}\biggl(v_{weak}+\frac{h(X)}{\sqrt{2}}\biggr){+h.c.}.$ (64)
IV.2 Integrating Out $r^{\mu}$ and Higher Dimension Operators
If we go beyond the pointlike limit in the BEH-Yukawa interaction we will have Lorentz invariant $O(1/M_{0}^{2})$ corrections. Before taking the pointlike limit of the Yukawa interaction, as in eq.83), we can rewrite this in the broken phase as:
$\!\!\!\!\!\!\!\!\!\!S_{Y}^{\prime}=m_{top}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{L}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}(M_{0}^{2}D_{F}(2r^{\mu}))\;F(r^{\mu})+h.c.,\qquad F=\frac{\widetilde{\Phi}(r)}{\widetilde{\Phi}(0)}\sim 1+a_{1}r^{2}+$... (65)
Note that,
$\!\!\!\!\!\!\!\!\!\!(D_{F}(2r^{\mu}))(r^{p})\sim\int\!\!\frac{d^{4}q}{(2\pi)^{4}}\frac{((\partial_{q})^{p}e^{iqr})}{q^{2}-M_{0}^{2}}\sim\frac{1}{M_{0}^{2+p}}\delta^{4}(r),$ (66)
implying that the expansion in $r$ corresponds to an expansion in $M_{0}^{-1}$, and derivatives $\partial_{X}^{p}$. The expansion in $r^{\mu}$ in all terms in the integrand generates a tower of assorted Lorentz invariant higher dimension operators such as:
$\frac{m_{t}}{M_{0}^{2}}\biggl(1+\frac{h(X)}{\sqrt{2}v}\biggr)\biggl([\overline{b}_{L}(X)D^{2}t_{R}(X)]+[\overline{t}_{L}(X)D^{2}t_{R}(X)]+...\biggr)+h.c.,$ (67)
$\overline{t}\rightarrow b+W+(g,\gamma,Z))\qquad\text{and,}\qquad\overline{t}\rightarrow t+(g,\gamma,Z)).$ (68)
$(\ell^{+}\ell^{-})\rightarrow(\gamma^{*},Z^{*})\rightarrow t+b+(W,g,\gamma,Z)).$ (69)